Determine the volume of the solid generated when the region bounded by the curve y^2=x^3 and the x-axis, is revolved about the line y=2

1
Expert's answer
2022-04-01T05:12:14-0400

Region bounded by the curve "y^2=x^3" and the x-axis, is revolved about the line "y=2".

Point of interception

"2^2=x^3" ; "x=4^{1\/3}"



Let's move to new coordinates:

"u=y-2\\t=x"

Now region is revolved about the line "u=0"

"(u+2)^2=t^3"

"u=t^{3\/2}-2"



The required volume can be obtained by subtracting volume A


"A=\int_0^{4^{1/3}} \pi(t^{3/2}-2)^2dt"

from volume B


"B=\int_0^{4^{1/3}} \pi(-2)^2dt=\int_0^{4^{1/3}} \pi2^2dt" .

"V=B-A=\int_0^{4^{1/3}} \pi(2^2-(t^{3/2}-2)^2)dt=""\int_0^{4^{1/3}} \pi(-t^3+4t^{3/2})dt=\pi(-\frac{t^4}{4}+\frac85t^{5/2})|_0^{4^{1/3}}=""\pi(-\frac{4^{4/3}}{4}+\frac854^{5/6})=\frac{\pi \cdot4^{1/3}\cdot 11}{5}\approx10.97"